the sum of 2 numbers is 54, and their difference is nine more than the smaller number.
Answer by checkley77(12569)

X+Y=54 OR X=54-Y
X-Y=X+9
X-X-Y=9
-Y=9
Y=-9 ANS.
X=54+9
X=63 ANS.
63-(-9)=63+9
72=72